OpenStudy (anonymous):

What is the limit of the sequence tan(2*n*pi/1+8*n). Thank You :)

ganeshie8 (ganeshie8):

Could you please format the question more better using latex / appropriate parenthesis ? :)

OpenStudy (anonymous):

\[a _{n}=\tan (\frac{ 2n \pi }{ 1+ 8n })\]

ganeshie8 (ganeshie8):

factor out "n" from the denominator, cancel it with the numerator "n" then take the limit

ganeshie8 (ganeshie8):

\[\large \lim\limits_{n\to\infty}a _{n}= \lim\limits_{n\to\infty}\tan \left(\frac{ 2n \pi }{ n\left[\frac{1}{n}+ 8\right]}\right)\]

ganeshie8 (ganeshie8):

\[\large = \lim\limits_{n\to\infty}\tan \left(\frac{ 2 \pi }{ \frac{1}{n}+ 8}\right)\]

ganeshie8 (ganeshie8):

take the limit